Department G had 1,800 units 25% completed at the beginning of the period, 12,200 units were completed during the period, 1,500 units were 20% completed at the end of the period, and the following manufacturing costs were debited to the departmental work in process account during the period:

Work in process, beginning of period $34,400
Costs added during period:
   Direct materials (11,900 units at $9) 107,100
   Direct labor 120,500
   Factory overhead 36,150

All direct materials are placed in process at the beginning of production, and the first-in, first-out method of inventory costing is used. What is the total cost of 1,800 units of beginning inventory that were completed during the period?

a.$48,440

b.$53,450

c.$51,950

d.$34,400
Answer #1

Equivalent unit of material = 11900

Equivalent unit of conversion = 1800*75%+10400+1500*20% = 12050

Cost per equivalent unit of material = 9

Cost per equivalent unit of conversion = 156650/12050 = 13

Total Cost of beginning inventory = 34400+1350*13 = 51950

So answer is c) $51950